OpenStudy (pottersheep):

Find "x" please. x + 2x^1/2 + 1 = 0?

jhonyy9 (jhonyy9):

\[x+2\sqrt{x}+1=0\]

OpenStudy (anonymous):

Substitute \[ u=\sqrt{x} \] You get \[ u^2 + 2u + 1 = (u+1)^2 = 0 \]

jhonyy9 (jhonyy9):

you want write this ?

OpenStudy (anonymous):

So u = -1, which means \[ -1=\sqrt{x} \] That proves that there is no real solution.